OpenStudy (loser66):

Prove that x^2+4 is an irreducible polynomial in R[x] Please help

OpenStudy (zzr0ck3r):

Do you have the Eisenstein criterion yet?

OpenStudy (loser66):

Not yet.

OpenStudy (zzr0ck3r):

hmm its easy with that, \(2|4\) but \(2^2\cancel{|}1\)

OpenStudy (zzr0ck3r):

What all do you have?

OpenStudy (loser66):

I stuck at this: Suppose x^2 +4 is reducible, then it can be factored as (ax+b)(cx+d) , it gives me x = -b/a and x =-d/c are roots of the x^2+4 but how to argue that they are not in real?

OpenStudy (zzr0ck3r):

Can you use the fact that this can have at most two roots?

OpenStudy (loser66):

another way to approach is open the bracket and get \[acx^2+x(ad+bc)+bd=x^2+4\] that is ac = 1 ad+bc =0 bd =4 then??

OpenStudy (zzr0ck3r):

Sec I am getting my book out

OpenStudy (loser66):

Back to the first option. if x =-b/a is a root, then (-b/a)^2 +4 =0 or \(\dfrac{b^2}{a^2}+4=0\) then??

OpenStudy (zzr0ck3r):

Ok do you know about the whole f(x) = g(x)h(x)+r(x) where r(x) has degree less than or equal to g(x)

OpenStudy (loser66):

yes.

OpenStudy (zzr0ck3r):

ok just trying to find out what all we can use

OpenStudy (anonymous):

\(\color{blue}{\text{Originally Posted by}}\) @Loser66 another way to approach is open the bracket and get \[acx^2+x(ad+bc)+bd=x^2+4\] that is ac = 1 ad+bc =0 bd =4 then?? \(\color{blue}{\text{End of Quote}}\) Well, this means that a and c have the same sign, so does b and d

OpenStudy (anonymous):

So consider the four cases.

OpenStudy (anonymous):

a and c are positive, b and d are positive.. then?

OpenStudy (loser66):

ac + bd \(\neq 0\)

OpenStudy (anonymous):

Then you can't have 0 in this case,

OpenStudy (zzr0ck3r):

Do you know that if it is degree 2 or 3, then it is irreducible over a field F if and only if it has a zero in F?

OpenStudy (zzr0ck3r):

This is not hard to prove

OpenStudy (zzr0ck3r):

then we can use the theorem of no more roots than n and we are done

OpenStudy (anonymous):

If a and c are positive and b and d are negative?

OpenStudy (loser66):

@wio then ac = bd

OpenStudy (anonymous):

Then ad is negative and bc is negative, so they can't add up to 0

OpenStudy (anonymous):

How about saying that \[ x^2 + 4 > 0 \] for any real x. So it cannot have real roots. So it cannot be factored over R. Am I missing something?

OpenStudy (anonymous):

@Loser66 Do you think you can do the rest now?

OpenStudy (loser66):

@zzr0ck3r If approach on that way, we have to prove why reducible polynomial has a root :)

OpenStudy (loser66):

@wio yes, Sir. It gives contradiction when ac =1 and bd =4 , right?

OpenStudy (loser66):

@eliassaab yes, it is quite clear. Can I do it??? I am doubt myself :)

OpenStudy (loser66):

Oh, at the end up, it has many ways to solve. Thanks you all. :)

OpenStudy (anonymous):

Oh, his way is simplest

OpenStudy (anonymous):

It just needs to be understood that you need real roots to be reducible

OpenStudy (loser66):

:)

OpenStudy (anonymous):

@Loser66 yes, you can it take it to the bank.

OpenStudy (zzr0ck3r):

@Loser66 C Is an extension of R, if we can show that there are two roots in C\R (very easy), then we are done since there are no roots in R, the degree 2 polynomial is irreducible.

OpenStudy (zzr0ck3r):

@eliassaab I think this is true iff degree 2 or 3 I have this proof in my notes \(f(x)\in F(x)\) then for \(f(x) \) degree \(2\) or \(3\) we have \(f(x)\) is irreducible if and only if \(f\) has a root in \(f\). I am not sure this is true for the other cases? Either way as I am sure they are starting from scratch, it should prob be proved.